OpenStudy (melissa_something):

A Math Riddle: (PLEASE HELP) One number is 6 less than a second number. Twice the second number is 42 more than 5 times the first. Find the two numbers.

OpenStudy (anonymous):

THANKS FOR PARTICIPATING IN THE 1ST MEDAL QUESTION

OpenStudy (anonymous):

I HAVE ONE PUT UP SO CHECK IT OUT BEFORE SOMEONE GETS THERE

OpenStudy (melissa_something):

Oh damn

OpenStudy (anonymous):

oh wow no one helps you

OpenStudy (melissa_something):

I knowwwww it's aggravating @KayleeNeedsHelp

OpenStudy (anonymous):

that has happened to a lot of people don't worry. I would help if I could

OpenStudy (anonymous):

can we do a medal exchange

OpenStudy (melissa_something):

Okay

OpenStudy (anonymous):

i gave you a medal now you give me one

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\text{let first number be x-6}\\~\\ \text{thus the second number would be x}\\~\\ 2x=42+5(x-6) }\end{align}\)

OpenStudy (melissa_something):

Youre amazing @mathmath333

OpenStudy (melissa_something):

So, distribute the 5 so it will be 2x=42+5x-30

OpenStudy (melissa_something):

Combine like terms? @mathmath333

OpenStudy (anonymous):

see, you found your help, never worry!

OpenStudy (mathmath333):

x should come -4 here